### v1

# Minimizing False-Positive Attributions in Explanations of Non-Linear Models

Abstract
Suppressor variables can influence model predictions without being dependent on the target outcome and they pose a significant challenge for Explainable AI (XAI) methods. These variables may cause false-positive feature attributions, undermining the utility of explanations. Although effective remedies exist for linear models, their extension to non-linear models and to instance-based explanations has remained limited. We introduce PatternLocal, a novel XAI technique that addresses this gap. PatternLocal begins with a locally linear surrogate, e.g. LIME, KernelSHAP, or gradient-based methods, and transforms the resulting discriminative model weights into a generative representation, thereby suppressing the influence of suppressor variables while preserving local fidelity. In extensive hyperparameter optimization on the XAI-TRIS benchmark, PatternLocal consistently outperformed other XAI methods and reduced false-positive attributions when explaining non-linear tasks, thereby enabling more reliable and actionable insights.
1 Introduction
In recent years, substantial advancements in explainable artificial intelligence (XAI) have been driven by the growing demand for transparent, accountable, and ethical AI systems across various industries and research communities. The need to understand machine learning models’ decision-making processes is now well-established, particularly in high-stakes domains such as healthcare, finance, and criminal justice. While promising progress has been made with so-called self-explainable models such as prototypical networks, these approaches often fall short of achieving the same downstream performance as conventional models and typically require complete retraining [18]. Consequently, significant attention remains focused on post-hoc XAI methods that can be applied to existing models without retraining.
Figure 1: Conceptual overview of how local linearization (top row) and PatternLocal (both rows) generate explanations. From the dataset (a), we train a classifier, whose decision boundaries are visualized using 2D projections. Traditional local XAI methods (top) approximate the model with a linear surrogate around the instance of interest. In contrast, PatternLocal (both rows) enhances this process by transforming the local discriminative surrogate into a generative explanation, effectively suppressing the impact of irrelevant or misleading features. The resulting importance maps (c, d) can be directly compared to the ground-truth attributions (b).
Post-hoc XAI methods encompass a diverse array of approaches, including architecture-specific techniques such as Layer-wise Relevance Propagation (LRP) [1], concept-based explanations like TCAV [13], and model-agnostic frameworks such as SHAP [15] and LIME [17]. These methods promise to support model validation, data quality checks, and actionable intervention suggestions [19]. To provide evidence, XAI methods are evaluated against criteria such as faithfulness to the model, robustness to perturbations, and human interpretability [26]. However, none of these criteria guarantee that the features highlighted as important are discriminative for the target variable, an assumption behind many uses of XAI [12].
Several studies [28, 29] have shown that widely used XAI methods often assign importance to suppressor variables [8, 27], features that affect model predictions without having a direct statistical dependency with the target. Suppressors may capture correlations in noise or other auxiliary signals that models exploit to improve predictions without offering causal or correlational insights. For instance, a model might predict epilepsy from seizure activity in a specific brain region but to optimally do so also rely on a noise probe from an unaffected region; common XAI methods may then highlight both regions, mistakenly assigning significance to the irrelevant area.
To support interpretations that rely on statistical relevance, a feature should be dependent on the target variable [28, 12]. The Pattern method [11] formalizes this idea by separating prediction from explanation: a discriminative model makes predictions, and a corresponding generative model provides explanations. Extensions such as PatternNet and PatternAttribution generalize this idea to deep neural networks [14] but have shown weaknesses on non-linear suppressor benchmarks like XAI-TRIS [7]. Unlike human-centered evaluations, benchmarks like XAI-TRIS provide controlled environments with objective ground-truth explanations to evaluate the robustness of XAI methods against suppressors and other uninformative variables. Also, model-agnostic methods such as SHAP, LIME, and LRP largely ignore training data structure, making them prone to suppressor bias [28, 29, 7].
In this work, we introduce PatternLocal, a data-driven, model-agnostic XAI method that extends the Pattern approach to non-linear cases. PatternLocal converts locally discriminative explanations, produced by methods such as LIME or KernelSHAP, into generative representations that more accurately reflect statistical relevance; see Figure 1. This transformation helps reduce the impact of suppressor and other non-informative variables while maintaining local fidelity to the model’s decision behavior. PatternLocal assumes access to representative training data and a meaningful input simplification (e.g. superpixels or feature masks) shared across samples. We evaluate PatternLocal on the XAI-TRIS suppressor and artificial lesion MRI benchmark and compare it with a range of established XAI methods. The results show that PatternLocal consistently provides more reliable and interpretable explanations without requiring any changes to the underlying model.
2 Preliminary
2.1 Notation
Let $\mathcal{X}\subset\mathbb{R}^{D}$ and $\mathcal{Y}\subset\mathbb{R}$ denote the input and output spaces, respectively. Consider a model $f:\mathcal{X}\rightarrow\mathcal{Y}$. For an instance $\mathbf{x_{\star}}\in\mathcal{X}$ we wish to explain the prediction $f(\mathbf{x_{\star}})$. Because highly complex models such as deep networks are notoriously hard to interpret, they are approximated locally by a simpler, intrinsically interpretable model $g$. Throughout we therefore restrict attention to local explanations for the specific instance $\mathbf{x_{\star}}$ rather than to the behavior of $f$ on $\mathcal{X}$ as a whole.
Many XAI methods first replace $\mathbf{x_{\star}}$ with a lower-dimensional or more semantically meaningful representation $\mathbf{x^{\prime}_{\star}}=h_{\mathbf{x_{\star}}}(\mathbf{x_{\star}})\in\mathbb{R}^{D^{\prime}}$. In image analysis, for example, $\mathbf{x^{\prime}_{\star}}$ may encode super-pixels; in other settings, it may correspond to a low-rank projection or simply a subset of raw features. Such simplification often makes for easier interpretation and, when $D^{\prime}\ll D$, also reduces the computational cost.
Remark.
In general, models that predict $y$ from $\mathbf{x}$ are termed backward or discriminative, whereas models that reconstruct $\mathbf{x}$ (or its sources) from latent factors are called forward or generative. We adopt this convention to avoid terminological clashes.
2.2 Suppressor Variables
Various studies [e.g., 8, 11] demonstrated that linear models may need to assign significant non-zero weight to correlated noise variables (suppressors). Given $\mathbf{x}\in\mathbb{R}^{D}$, assume each instance is generated from a linear model of the form
$\mathbf{x}\;=\;\mathbf{A}\,\mathbf{s}\;+\;\bm{\varepsilon},\qquad\mathbf{s}=\bigl{[}s_{1},\dots,s_{K}\bigr{]}^{\!\top}\in\mathbb{R}^{K},$ (1)
where $K$ is the number of latent factors, $\mathbf{A}=[\mathbf{a}_{1},\dots,\mathbf{a}_{K}]\in\mathbb{R}^{D\times K}$ contains the activation patterns $\mathbf{a}_{k}$, and $\bm{\varepsilon}$ denotes additive noise. A linear backward model attempts to recover the latent vector $\mathbf{s}$ via
$\hat{\mathbf{s}}\;=\;\mathbf{W}^{\!\top}\mathbf{x},\qquad\mathbf{W}\in\mathbb{R}^{D\times K}.$ (2)
Consider $D=2$ and $K=1$ with $x_{1}=s+d$ and $x_{2}=d$, where the scalar $s$ is the signal of interest and $d$ is a suppressor. The signal is exactly recovered by $\hat{s}=x_{1}-x_{2}=\mathbf{w}^{\!\top}\mathbf{x}$ with $\mathbf{w}=[1,-1]^{\!\top}$. Interpreting $\mathbf{w}$ directly as feature importance, however, suggests $x_{1}$ and $x_{2}$ contribute equally, even though the second channel carries only the suppressor, which is removed entirely in $\hat{s}$ by subtraction. This issue with suppressor variables extends to common explainable AI methods used to interpret machine learning models [29, 7].
2.3 Forward vs. backward models
Instead of using the weights of the linear backward model, Haufe et al. [11] propose to use the corresponding forward model. They show that the weight matrix $\mathbf{W}$ has a unique forward-model counterpart
$\mathbf{A}\;=\;\Sigma_{\mathbf{X}}\,\mathbf{W}\,\Sigma_{\mathbf{S}}^{-1},$ (3)
where $\Sigma_{\mathbf{X}}=\operatorname{Cov}[\mathbf{x}]$ and $\Sigma_{\mathbf{S}}=\operatorname{Cov}[\hat{\mathbf{s}}]$ which recovers the activation pattern in Equation 1. If the estimated factors are uncorrelated, $\Sigma_{\mathbf{S}}$ is diagonal and Equation 3 reduces to $\mathbf{A}\;\propto\;\Sigma_{\mathbf{X}}\,\mathbf{W}\;=\;\operatorname{Cov}\bigl{[}\mathbf{x},\,\hat{\mathbf{s}}\bigr{]}$. The resulting global feature importance map for the entire data set $\mathcal{D}\subset\mathcal{X}$ is
$\mathbf{s}^{\text{Pattern}}(\mathcal{D})\;=\;\Sigma_{\mathbf{X}}\,\mathbf{W},$ (4)
which mitigates suppressor variables for the linear model. For example, in unregularized linear discriminant analysis, where $\mathbf{w}=\Sigma_{\mathbf{X}}^{-1}(\bm{\mu}_{+}-\bm{\mu}_{-})$ with $\bm{\mu}_{+/-}$ denoting the class means, this approach completely removes the influence of suppressor variables in the activation pattern.
2.4 Local explainability methods
LIME.
LIME works by constructing a local surrogate model $g$ of $f$ around $\mathbf{x^{\prime}_{\star}}$; the surrogate’s coefficients form the feature-importance explanation for $\mathbf{x}$ [17]. Its original formulation is $\xi\;=\;\arg\!\min_{g\in G}\;\mathcal{L}\!\left(f,g,\Pi_{\mathbf{x^{\prime}_{\star}}}\right)\;+\;\Omega(g)$, where $G$ is a class of interpretable functions, $\mathcal{L}$ is a loss evaluated on samples in the simplified space, $\Pi_{\mathbf{x^{\prime}_{\star}}}$ is a local kernel, and $\Omega$ penalizes model complexity. In the common usage where $\mathcal{L}$ is squared loss and $g$ is linear, can be written as
$\mathbf{s}^{\text{LIME}}(\mathbf{x}_{\star})\;=\;\mathbf{w}^{\text{LIME}}\;=\;\arg\!\min_{\mathbf{v}}\;\mathbb{E}_{\mathbf{z}^{\prime}\,\sim\,\mathcal{Z}}\Bigl{[}\Pi_{\mathbf{x^{\prime}_{\star}}}\!\bigl{(}\mathbf{z}^{\prime}\bigr{)}\bigl{(}f\!\bigl{(}h_{\mathbf{x_{\star}}}^{-1}(\mathbf{z}^{\prime})\bigr{)}-\mathbf{v}^{\!\top}\mathbf{z}^{\prime}\bigr{)}^{2}\Bigr{]}\;+\;\lambda\,R(\mathbf{v}),$ (5)
where $R$ is a regularization weighted by $\lambda$, $\mathcal{Z}$ is the simplified input space, and $h_{\mathbf{x_{\star}}}^{-1}\!:\mathcal{Z}\!\rightarrow\!\mathcal{X}$ maps a simplified sample back to the original space. This notation is similar to the one by Tan et al. [25]. For image explanations the default choices are $\Pi_{\mathbf{x^{\prime}_{\star}}}(\mathbf{z}^{\prime})=\exp\,\!(-\|\mathbf{1}-\mathbf{z}^{\prime}\|_{0}^{2}/\sigma^{2})$, $R(\mathbf{v})=\|\mathbf{v}\|_{2}^{2}$, and $h_{\mathbf{x^{\prime}_{\star}}}^{-1}(\mathbf{z}^{\prime})=\mathbf{z}=\mathbf{x_{\star}}\odot B(\mathbf{z}^{\prime})+\mathbf{r}\odot(\mathbf{1}-B(\mathbf{z}^{\prime}))$, with $\mathbf{z}^{\prime}\sim\text{Uni}(\{0,1\}^{D^{\prime}})$, $\odot$ denoting element-wise product, $r$ being a reference value, and where $B:\{0,1\}^{d}\rightarrow\{0,1\}^{D}$ copies each bit $z^{\prime}_{j}$ to every pixel in super-pixel. As mentioned, LIME can assign non-zero importance to suppressor variables [29].
KernelSHAP & Gradient methods.
As shown by Lundberg and Lee [15], Tan et al. [25], KernelSHAP can be written in the form of Equation 5 by setting $R(\mathbf{v})=0$ and $\Pi_{x^{\prime}}(z^{\prime})=(D^{\prime}-1)\big{/}\left[\left(D^{\prime}\,\mathrm{choose}\,\|z^{\prime}\|_{0}\right)\cdot\|z^{\prime}\|_{0}\cdot(D^{\prime}-\|z^{\prime}\|_{0})\right]$. Tan et al. [25] further demonstrate that Equation 5 unifies SmoothGrad [22] and the plain Gradient method [30].
3 PatternLocal
We now unify the Pattern insight that forward model parameters reveal statistically relevant features with the ability to probe highly non-linear models locally with methods like LIME. We name this PatternLocal. PatternLocal converts any linear surrogate explanation $\mathbf{w}$ produced around an instance $\mathbf{x}_{\star}$ into a data-driven pattern $\mathbf{a}$ that (i) suppresses false positives caused by suppressor variables, (ii) remains faithful to the model because it is built on the very surrogate that approximates the decision boundary, (iii) preserves local fidelity by weighting the surrounding data with a local kernel, and (iv) remains practical by operating in a simplified, low-dimensional space with regularized estimation.
(1) Suppressor-variable mitigation.
Pattern mitigates suppressor variables in linear models by converting the backward model into a forward one. This can be interpreted as regressing $\mathbf{x}$ on the estimated signal $\tilde{\mathbf{s}}=\mathbf{W}^{\top}\mathbf{x}$ under a Gaussian noise model. The goal is to reconstruct $\mathbf{x}$ from $\tilde{\mathbf{s}}$, effectively down-weighting features that do not carry the signal. PatternLocal extends this idea to the local setting by regressing the local surrogate output $\tilde{y}=\mathbf{w}^{\top}\mathbf{x}$ onto the neighborhood of $\mathbf{x}_{\star}$, thus removing suppressor effects specific to the instance.
(2) Model faithfulness.
PatternLocal inherits its surrogate $\mathbf{w}$ from Equation 5, which is explicitly optimized to approximate the model $f$ in the local neighborhood of $\mathbf{x}_{\star}$. By regressing the simplified input $\mathbf{x}$ on the surrogate model’s response $\tilde{y}=\mathbf{w}^{\top}\mathbf{x}$, PatternLocal preserves the local approximation. Crucially, we regress onto the surrogate output rather than the true label $y$, since $y$ reflects only the ground truth and not the behavior of the model $f$. Using $y$ instead would explain the task, not the model $f$, violating model faithfulness.
(3) Local fidelity.
In non-linear models, the important features can vary across different regions of the input space. PatternLocal enforces locality by weighting all terms in its regression objective using the local kernel $\Pi_{\mathbf{x^{\prime}_{\star}}}$. This ensures that the explanation focuses on the neighborhood around $\mathbf{x}_{\star}$ and does not reflect global structure.
(4) Feasibility and dimensionality.
PatternLocal requires access to the training data $\mathcal{D}$ and assumes that a sufficient number of samples lie in the neighborhood of $\mathbf{x}_{\star}$. To make this feasible, it operates in the same simplified space defined by $\mathbf{h}_{\mathbf{x_{\star}}}$ as in Equation 5, which not only aligns the explanation with the surrogate but also drastically reduces the dimensionality and the number of samples required. Nevertheless, the regression can still become ill-posed in sparse or high-dimensional settings. To ensure stability, we include a regularization $Q$.
3.1 Formal objective
Let $\mathbf{h}_{\mathbf{x_{\star}}}:\mathcal{X}\rightarrow\mathbb{R}^{D^{\prime}}$ be the simplified input representation used in Equation 5 and $\mathbf{w}$ be its solution. Also let $\tilde{y}=\mathbf{w}^{\!\top}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})$ denote the local surrogate prediction. PatternLocal estimates a pattern $\mathbf{a}$ by regressing the local prediction onto the simplified input of the original data within the local neighborhood of $\mathbf{x}_{\star}$
$\mathbf{s}^{\text{PatternLocal}}(\mathbf{x_{\star}})\;=\;\mathbf{a}\;=\;\arg\!\min_{\mathbf{u}}\;\mathbb{E}_{\mathbf{x}\,\sim\,\mathcal{X}}\Bigl{[}\Pi_{\mathbf{x^{\prime}_{\star}}}(\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x}))\,\bigl{(}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})-\mathbf{u}\,{\tilde{y}}\bigr{)}^{2}\Bigr{]}\;+\;\lambda\,Q(\mathbf{u}).$ (6)
When $Q(\mathbf{u})=\|\mathbf{u}\|_{1}$, the objective can be solved with a kernel-weighted Lasso regression. For $Q(\mathbf{u})=\|\mathbf{u}\|_{2}^{2}$ the problem reduces to a kernel-weighted ridge regression with a closed-form solution given by
$\mathbf{a}_{\ell_{2}}\;=\;\bigl{(}\,\mathbb{E}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\big{[}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})\,\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})^{\!\top}\bigr{]}+\lambda I\bigr{)}^{-1}\,\mathbb{E}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\big{[}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x})\,\tilde{y}\,\big{]}\;=\;\frac{\operatorname{Cov}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\!\bigl{[}\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x}),\,\tilde{y}\bigr{]}}{\operatorname{Var}_{\Pi_{\mathbf{x^{\prime}_{\star}}}}\!\bigl{[}\tilde{y}\bigr{]}\;+\;\lambda},$ (7)
where expectations, covariances, and variances are taken with respect to the normalized kernel $\Pi_{\mathbf{x^{\prime}_{\star}}}(\mathbf{h}_{\mathbf{x_{\star}}}(\mathbf{x}))$ over $\mathcal{X}$. See supplementary material for derivations. Equation 7 therefore expresses $\mathbf{a}_{\ell_{2}}$ as the kernel-weighted covariance between the simplified features and the surrogate response, scaled by its regularized variance.
Figure 2: Feature–importance comparison on the XOR toy problem. We draw 2 50025002\,5002 500 i.i.d. samples from the generative process of Equation 8, so the label depends solely on the interaction between $x_{1}$ and $x_{2}$, while $x_{3}$ is a suppressor that carries no marginal predictive signal. Each plot shows $(x_{1},x_{2})$ pairs, colored by the local normalized feature importance that four XAI methods assign to the suppressor variable $x_{3}$ when applied to the smooth classifier. Every plot reports the empirical mean magnitude of this attribution across all test points. Whereas LIME, KernelSHAP, and Gradient all attribute substantial importance to $x_{3}$, PatternLocal correctly drives the attribution of $x_{3}$ to (almost) zero.
3.2 Toy example
To better illustrate the intuition behind the PatternLocal method, we present a small-scale toy example. Let $s_{1},s_{2}\sim\mathcal{N}(0,1)$ and $d\sim\mathcal{N}(0,\sigma_{d}^{2})$ be independent latent variables. We observe
$\mathbf{x}=\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\end{bmatrix}=\begin{bmatrix}s_{1}+d\\ s_{2}-d\\ d\end{bmatrix},\qquad y=\operatorname{sign}(s_{1}s_{2}),$ (8)
which induces a classical XOR-like dependency between $(x_{1},x_{2})$ and the label $y$. The variable $x_{3}$ acts as a suppressor. The optimal classifier for this non-linear problem is given by $\tilde{y}=\operatorname{sign}(\tilde{f}(\mathbf{x}))$, $\tilde{f}(\mathbf{x})=(x_{1}-x_{3})(x_{2}+x_{3})$. A smooth approximation to this classifier is $f_{\tau}(\mathbf{x})=\tanh((1/\tau)\,\tilde{f}(\mathbf{x}))$, $0<\tau\ll 1$. Fixing an input $\mathbf{x}_{\star}$, the first-order (linear) approximation of the decision function at this point is $f_{\tau}(\mathbf{x})\approx f_{\tau}(\mathbf{x}_{\star})+\mathbf{w}^{\top}(\mathbf{x}-\mathbf{x}_{\star})$ with
$\mathbf{w}=\nabla\tilde{f}(\mathbf{x}_{\star})=c_{\star}\begin{bmatrix}{x_{2}}_{\star}+{x_{3}}_{\star}\\ {x_{1}}_{\star}-{x_{3}}_{\star}\\ {x_{1}}_{\star}-{x_{2}}_{\star}-2{x_{3}}_{\star}\end{bmatrix},\qquad c_{\star}=\frac{1}{\tau}\operatorname{sech}^{2}\left(\frac{1}{\tau}\tilde{f}(\mathbf{x}_{\star})\right),$ (9)
noting that $w_{1}-w_{2}+w_{3}=0$. Therefore, we in general have $w_{3}\neq 0$, implying that local linear XAI methods (e.g., gradients, LIME, KernelSHAP) will incorrectly assign nonzero importance to the suppressor variable $x_{3}$.
Let $\Pi_{\mathbf{x}_{\star}}(\mathbf{x})=\varphi(\|\mathbf{x}-\mathbf{x}_{\star}\|)$ be any isotropic local kernel. For $\lambda=0$ and identity mapping $h_{\mathbf{x}_{\star}}(\mathbf{x})=\mathbf{x}$, the PatternLocal explanation vector is defined as $\mathbf{a}=\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[\mathbf{x},\tilde{y}]\big{/}\operatorname{Var}_{\Pi_{\mathbf{x}_{\star}}}[\tilde{y}]$ and $\tilde{y}=\mathbf{w}^{\top}\mathbf{x}$.
$\mathbf{a}=\frac{\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[\mathbf{x},\tilde{y}]}{\operatorname{Var}_{\Pi_{\mathbf{x}_{\star}}}[\tilde{y}]},\qquad\tilde{y}=\mathbf{w}^{\top}\mathbf{x}.$
Consider the third component $a_{3}=\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[x_{3},\tilde{y}]/\operatorname{Var}_{\Pi_{\mathbf{x}_{\star}}}[\tilde{y}]$, we compute
$\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[x_{3},x_{1}]=\sigma_{d}^{2},\qquad\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[x_{3},x_{2}]=-\sigma_{d}^{2},\qquad\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[x_{3},x_{3}]=\sigma_{d}^{2}.$
Thus,
$\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[x_{3},\tilde{y}]$ $=w_{1}\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[x_{3},x_{1}]+w_{2}\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[x_{3},x_{2}]+w_{3}\operatorname{Cov}_{\Pi_{\mathbf{x}_{\star}}}[x_{3},x_{3}]$
$=w_{1}\sigma_{d}^{2}-w_{2}\sigma_{d}^{2}+w_{3}\sigma_{d}^{2}=\sigma_{d}^{2}(w_{1}-w_{2}+w_{3})=0,$
and hence $a_{3}=0$. Similar calculation can be done for $a_{1}$ and $a_{2}$ which yields $\mathbf{a}\propto[w_{1},w_{2},0]^{\top}$. See supplementary material for derivations. We conclude that PatternLocal correctly assigns zero feature importance to the suppressor variable $x_{3}$, in contrast to standard local linear methods. This is also confirmed with simulation as seen in Figure 2.
4 Experiments
The experimental workflow begins with the construction of synthetic image datasets whose pixel-level ground-truth attribution maps are known by design, giving us full control over class-conditional distributions [^1] Code is available at https://github.com/gjoelbye/PatternLocal. Each candidate classifier is trained on a standard train/validation split and its performance is subsequently assessed on an unseen test set. For every test image, the selected XAI method produces a normalized importance map in the range $[-1,1]$. These maps are then quantitatively compared with the ground-truth attributions using the evaluation metrics introduced below. All XAI hyperparameters are tuned on the validation split before the final assessment to ensure a fair comparison.
4.1 Datasets
We focus on the XAI-TRIS benchmark dataset [7] and the artificial lesion MRI dataset [16]. Both datasets are specifically designed to assess XAI methods in the context of suppressor variables.
XAI-TRIS Benchmark
We follow the methodology outlined in Clark et al. [7]. Full details are provided in supplementary materials. The XAI-TRIS benchmark dataset consists of $64\times 64$ images (with $N=40{,}000$ samples) generated by combining a class-dependent foreground tetromino shape with background noise. The relative weighting between signal and noise is set by $\alpha$. Two noise scenarios are considered:
  • WHITE: uncorrelated Gaussian noise,
  • CORR: spatially smoothed Gaussian noise (leading to suppressor variables).
We evaluate three scenarios:
  • LIN: A ‘T’-shaped tetromino near the top-left indicates $y=0$; an ‘L’-shaped tetromino near the bottom-right indicates $y=1$.
  • XOR: Both tetrominoes appear with differing signs depending on $y$.
  • RIGID: Each tetromino is randomly translated and rotated by 90 increments.
The set of important (ground-truth) pixels is determined by the tetromino shape. For training and evaluation, each dataset is split into $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{val}}$, and $\mathcal{D}_{\text{test}}$ in a 90/5/5 ratio. Figure 3 shows examples of all scenarios. Spatially correlated noise in the CORR setting introduces suppressor pixels, background locations correlated with the foreground through the smoothing operator $G$. Although these pixels carry no direct class information, the model can exploit them to denoise the true signal [11, 28].
(a) Scenario CORR with $\alpha=0.2$, where noise is correlated with the label, introducing a suppressor variable.
(b) Scenario WHITE with $\alpha=0.2$, where noise is independent (white), and no suppressor variable is present.
Figure 3: Examples of $64\times 64$ instances from the XAI-TRIS Benchmark across six different scenario types. Each row represents a distinct structural pattern (LIN, XOR, RIGID), and each column pair shows a binary label class ($y=0$ and $y=1$) along with the corresponding instance and attribution mask. On the right (3(a)), correlated noise introduces a spurious suppressor effect, while on the left (3(b)), white noise does not.
Artificial lesion MRI dataset
[…]
4.2 Models
We consider two types of classifiers; a multi-layer perceptron (MLP) and a convolutional neural network (CNN). Each classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$ is trained on the training dataset $\mathcal{D}_{\text{train}}$. We train the model with early stopping based on the validation performance on $\mathcal{D}_{\text{val}}$ and then we evaluate the model of the test set $\mathcal{D}_{\text{test}}$. We consider a classifier to have generalized the given classification problem when the resulting test accuracy is above 90%. See supplementary material for details on model architecture, training, and performance.
4.3 XAI methods
We evaluate 11 methods on the models, including the PatternLocal approach and a baseline. The interpretability methods applied are: LIME [17], Integrated Gradients [24], Saliency [21], DeepLift [20], GradientShap [9], and GuidedBackProp [23]. Additionally, we incorporate two edge detection filters, Sobel and Laplace, which have been shown to outperform more advanced XAI techniques in suppressor variable scenarios [7]. Note that KernelSHAP is a special case of LIME. Since we perform extensive hyperparameter optimization, KernelSHAP may also be selected during this process.
For the XAI-TRIS dataset we benchmark LIME and PatternLocal with three input-simplification schemes for $\mathbf{h}_{\mathbf{x}_{\star}}(\cdot)$: the identity mapping, a super-pixel representation, and a low-rank approximation. The same three variants are evaluated on the artificial-lesion MRI dataset. Super-pixels are generated on a uniform grid for XAI-TRIS, whereas the slic algorithm is used for the MRI data.
4.4 Metrics
Earth Mover’s Distance (EMD).
The Earth Mover’s Distance (EMD) measures the optimal cost required to transform one distribution into another. For a continuous-valued importance map $s$ and ground truth $F^{+}$, both normalized to have the same mass, the EMD is computed using the Euclidean distance as the ground metric. The optimal transport cost, $\text{OT}(|\mathbf{s}|,F^{+})$, is calculated following the algorithm proposed by Bonneel et al. [6] and implemented in the Python Optimal Transport library [10]. We define a normalized EMD performance score as
$\text{EMD}={\text{OT}(|\mathbf{s}|,\mathcal{F}^{+})}/{\delta_{\textit{max}}}\;,$ (10)
where $\delta_{\textit{max}}$ is the maximum Euclidean distance between any two pixels.
Importance mass error (IME).
As argued in Clark et al. [7] and Arras et al. [2], it is plausible that the given model only uses a subset of the important pixels for its prediction. Thus, false positives should be preferentially penalized while false negatives should be largely ignored. To this end, we provide an additional metric, Importance Mass Error (IME),
$\text{IME}=1-\sum_{i=1}^{|\mathcal{F}^{+}|}|s_{i}|/\sum_{i=1}^{|\mathbf{s}|}|s_{i}|.$ (11)
4.5 Hyperparameter optimization
To ensure a fair evaluation of the XAI methods, we perform computationally intensive hyperparameter optimization for each uniquely generated dataset, scenario, and signal-to-noise ratio (SNR), as well as for the corresponding trained machine learning models. We use Bayesian optimization with the Tree of Parzen Estimators (TPE) algorithm [4], which is well-suited for handling the mix of continuous and categorical hyperparameters in our setup. This optimization is carried out using the hyperopt Python package [5]. The loss function we minimize is the Earth Mover’s Distance (EMD) between the absolute value of the normalized explanation and the ground truth. A detailed description of the hyperparameter selection process for the XAI methods is provided in the supplementary material.
5 Results
XAI-TRIS Benchmark
(c) Earth mover’s distance (EMD)
(d) Importance mass error (IME)
Figure 4: Quantitative evaluation of feature importance maps for the MLP model using $64\times 64$ XAI-TRIS benchmark images generated by various methods, as a function of the SNR. The MLP model achieves at least $90\%$ accuracy, and method hyperparameters were optimized for EMD. The function $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ is the identity function, and $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$. The top panel (4) shows results based on EMD, while the bottom panel (4) presents performance in terms of IME. PatternLocal demonstrates significant improvements over most methods. Shaded regions indicate standard deviation as the standard error is too small to be visible.
We evaluate the performance of the XAI methods on the XAI-TRIS benchmark both qualitatively and quantitatively. Qualitative evaluation is done through visual inspection, which is standard in XAI research [e.g., 3, 15, 7], and illustrated in supplementary material. The quantitative evaluation uses ground truth along with the EMD and IME metrics, as shown in Figure 4. The results presented here are for the MLP model, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ as the identity function and $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$. Results for the CNN model, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$ as superpixels or low-rank approximation, and for $R(\cdot)=Q(\cdot)=\|\cdot\|_{1}$, are provided in the supplementary material.
Figure 5: Examples from the artificial lesion MRI dataset, showing two MRI slices (rows). The first column displays the original MRI slices with artificially generated lesions; the second column (Ground Truth) shows the true lesion locations. The subsequent columns show feature importance maps from PatternLocal, LIME, Sobel, and Laplace. We use a CNN model, $\mathbf{h}_{\mathbf{x_{\star}}}(\cdot)$, with superpixels, and set $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$ for both PatternLocal and LIME. Note that the filter-based methods fail in these examples due to the lack of clear edges. Overlaid on the ground truth are superpixels with feature importance above 0.9 from PatternLocal (green) and LIME (red). PatternLocal explanations appear better aligned with the lesion locations. The overlap is not perfect, as it is limited by the superpixel resolution.
In Figure 4, we present the aggregated results of the methods across the test set for the EMD and IME metrics. Overall, the PatternLocal method significantly outperforms all other methods in the XOR and RIGID scenarios. However, in the RIGID CORR scenario, the results are mixed: PatternLocal performs substantially better than other XAI methods but is comparable to or worse than the filter methods Sobel and Laplace. These filter methods also performed well in Clark et al. [7], which is due to the XAI-TRIS methods involving rigid bodies with sharp lines, making them especially easy for filter-based methods to detect. This advantage does not carry over to images with more complex backgrounds, such as those in the artificial lesion MRI dataset.
Additional ablation studies on the impact of hyperparameters and the role of the suppressor variable are provided in the supplementary material.
Artificial lesion MRI dataset
We evaluate the XAI methods using the same procedure on the artificial lesion MRI dataset. The results reported here correspond to the CNN model, with $\mathbf{h}_{\mathbf{x}_{\star}}(\cdot)$ defined over superpixels and $R(\cdot)=Q(\cdot)=\|\cdot\|_{2}^{2}$. Figure 5 illustrates the performance of PatternLocal in comparison to LIME and the filter-based methods Sobel and Laplace. Additional qualitative and quantitative benchmark results are provided in the supplementary material.
6 Conclusion
This paper presents PatternLocal, a data-driven model-agnostic XAI method that improves local explanations by reducing false-positive attributions caused by suppressor variables. By transforming standard local surrogate outputs into more reliable feature attributions, PatternLocal enhances interpretability without retraining the underlying model. Across two benchmarks, XAI-TRIS and an artificial lesion MRI dataset, PatternLocal consistently outperforms existing methods, offering clearer, more trustworthy explanations, especially in challenging scenarios with misleading noise.
Limitations
The primary limitation of PatternLocal is its reliance on access to representative training data in the neighborhood of each explanation. While this is feasible, it may be challenging in privacy-sensitive or data-scarce environments. Additionally, because PatternLocal fits a local generative model, it assumes a degree of consistency or alignment across samples in the input space. This assumption holds well in structured domains like medical imaging, where standardized views are common, but may break down in less constrained settings, such as natural images or user-generated content, where spatial alignment across instances cannot be guaranteed. PatternLocal reduces false-positive feature attribution, however like other XAI methods it can still misattribute importance, so saliency maps should be seen as suggestive rather than definitive.
Acknowledgments
We thank Sebastian Weichwald, William Lehn-Schiøler, Teresa Karen Scheidt, and Lina Skerath for their contributions during the early stages of this work, particularly through their involvement in the Master’s thesis that inspired it. We are also grateful to the anonymous reviewers for their valuable feedback. This work was supported by the Pioneer Centre for AI (DNRF P1) and the Novo Nordisk Foundation (“Cognitive spaces – Next generation explainability”, NNF22OC0076907), as well as the Danish Data Science Academy, funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516).